OpenStudy (anonymous):

Solve the systems of equations using elimination. 8x+7y=1 and 5x-4y=-3 HELP test tomorrow and dont understand. Thanks

jhonyy9 (jhonyy9):

8x+7y=1 multiplie by 5 5x-4y=-3 multplie by -8 40x+35y=5 -40x+32y=24 --------------- 0 67y=29 y = 29/67 hope you can continue it for x sure